### v1

@@ Abstract
Algorithms for computing classical Gaussian quadrature rules (Gauss-Jacobi, Gauss-Laguerre, and Gauss-Hermite) are presented, based on globally convergent fourth-order iterative methods and asymptotic approximations, which are applied in complementary regions of the parameter space. The combination of these approaches results in methods that surpass previous algorithms in terms of speed, accuracy, and computational range (practically unrestricted). The Gauss-Radau and Gauss-Lobatto variants are also considered, along with the computation of the associated barycentric weights. Arbitrary accuracy algorithms are also provided for the symmetric cases (Gauss-Gegenbauer and Gauss-Hermite).
@@ 1 Introduction
where $w(x)$ is a weight function defined on the interval $[a,b]$ (with $a$ and/or $b$ possibly infinite), are quadrature rules with maximal polynomial exactness. Specifically, they satisfy $I(x^{k})=Q_{n}(x^{k})$ for $k=0,1,\dots,2n-1$. These rules exhibit fast convergence, particularly for analytic functions, making them widely used in numerous applications, including spectral methods.
As is well known, the nodes $x_{i}$ of an $n$-point Gaussian quadrature rule correspond to the zeros of the orthogonal polynomial $p_{n}(x)$ of degree $n$ (unique up to a normalization factor) that satisfies $I(x^{k}p_{n}(x))=0,\quad k=0,1,\dots,n-1$. The quadrature weights are related to the values of the orthogonal polynomials at these nodes. The nodes and weights can also be expressed in terms of the eigenvalues and eigenvectors of the Jacobi matrix associated with the three-term recurrence relation satisfied by the family of orthogonal polynomials $\{p_{k}(x)\}_{k\in\mathbb{N}}$. See, for instance, [8, sect. 5.3].
Particularly significant are the so-called classical quadrature rules, for which explicit expressions for the orthogonal polynomials are known. These include Gauss-Hermite, Gauss-Laguerre, and Gauss-Jacobi quadratures. In these cases, the weights of the $n$-point rule can be expressed in terms of the derivative of the corresponding orthogonal polynomial evaluated at the nodes. Table 1 summarizes the fundamental relations for these quadratures.
@@
Table 1: Classical Gaussian quadratures: interval of integration $(a,b)$, weight function $w(x)$, orthogonal polynomial $p_{n}(x)$ and its leading order coefficient $k_{n}$ [16, table 18.3.1], and expression for the weights[9]. The constant $M_{n,\alpha,\beta}$ is $M_{n,\alpha,\beta}=2^{\alpha+\beta+1}\Gamma(n+\alpha+1)\Gamma(n+\beta+1)/(n!\Gamma(n+\alpha+\beta+1))$.
Numerical algorithms are presented for computing the three classical Gaussian quadrature rules, based on the asymptotic methods described in [9] (Gauss-Hermite and Gauss-Laguerre) and [11] (Gauss-Jacobi), as well as on the globally convergent iterative methods of [10] (Gauss-Hermite and Gauss-Laguerre) and [13] (Gauss-Jacobi). Asymptotic and iterative algorithms have distinct advantages and their combination, as we prove in this paper, yields the fastest and most accurate methods available, with the broadest range of validity with respect to the parameters.
Once efficient methods for computing classical Gaussian quadratures are available, related problems can also be addressed. These include computing modifications such as Gauss-Radau and Gauss-Lobatto rules, as well as evaluating barycentric weights for interpolation at the nodes of classical orthogonal polynomials.
@@
The paper is structured as follows. First, for each of the three classical Gaussian quadratures, we describe the different computational methods (iterative and asymptotic), emphasizing aspects specific to each approach that were not fully detailed in previous publications. We begin with Gauss-Jacobi quadrature, with Gauss-Chebyshev and Gauss-Legendre as particularly important cases. For the Legendre case, a highly efficient method based purely on asymptotics and a look-up table for low degrees is presented. After discussing Gauss-Jacobi quadrature, the Gauss-Laguerre and Gauss-Hermite cases are described. Subsequently, we discuss the computation of Gauss-Radau and Gauss-Lobatto quadratures and the evaluation of barycentric weights for interpolation at orthogonal polynomial nodes. Finally, the last section summarizes the set of Matlab and Maple algorithms accompanying this work.
2 Gauss-Jacobi quadrature
@@
We consider separately the Chebyshev cases $|\alpha|=|\beta|=1/2$, the Legendre case $\alpha=\beta=0$ and the more general Jacobi case (including Gauss-Gegenbauer quadrature $\alpha=\beta$). Before describing these cases in detail, we give a brief summary of some properties of the general Gauss-Jacobi case that will be useful in the subsequent discussion.
@@
The values $|\dot{Y}(\theta_{i})|$ have a small variation as a function of $i$. This is a consequence of the fact that the coefficient of the ODE is nearly constant as $n\rightarrow+\infty$ in any compact subinterval of (1,1)(-1,1). This is consistent with the known asymptotic approximations as $n\rightarrow+\infty$ (see for instance [19], Eq. (A.8)):
@@
2.1 Gauss-Chebyshev quadratures
For the Chebyshev cases, $|\alpha|=|\beta|=1/2$, simple explicit formulas exist for the nodes and weights [18, Thm. 8.4]. The four cases can be summarized in a single formula as follows:
@@ 2.1 Gauss-Chebyshev quadratures
Care must be taken in order to avoid loss of accuracy in the evaluation of the trignometric functions involved in (6). For instance, the previous expression for $x_{k}$ is not convenient for computing the nodes close to zero which, if $\alpha<<n$, correspond to values $k\simeq n/2$. Indeed, we have $x_{k}=\cos\theta_{k}$ with $\theta_{k}=\pi/2+\delta_{k}$, with $|\delta_{k}|$ small; then $x_{k}=-\sin(\delta_{k})\simeq\delta_{k}$ and $\delta_{k}=\theta_{k}-\pi/2$ suffers loss of accuracy by cancellation. In this case it is preferable to use the expression $x_{k}=-\sin((\frac{n}{2}-k+\frac{\alpha-\beta+2}{4})\frac{\displaystyle{\pi}}{\displaystyle{\kappa}}).$ where the small number n/2kn/2-k (compared to $\kappa$) is computed exactly. Similarly precautions must be taken for the factors $1\pm x_{k}$ appearing in the weights.
@@
2.2 Gauss-Legendre quadrature
The Gauss-Legendre case is a particularly important one and, in fact, it is costumary to use the expression “Gauss quadrature” to refer to this method alone. Because $x_{k}=-x_{n-k+1}$ ($1<x_{1}<...<x_{n}<1$) and $w_{k}=w_{n-k+1}$ we only need to compute the non-negative nodes and its corresponding weights.
For Gauss-Legendre quadrature, the Bessel-type expansions of Legendre polynomials [23] enable explicit asymptotic expressions for nodes and weights, as in [1]. The chebfun package legpts [3] follows this approach for $n>100$. We also use Bogaert’s expansions, but only for a fixed number of nodes and weights: for $n>80$, the 35 largest positive nodes and corresponding weights $w_{i}$ are computed using Bogaert’s formulas, while expansions in elementary functions are considered for the rest. For $n\leq 80$, we use a look-up table storing precomputed nodes and weights with full double precision ($16$ significant digits), generated using the arbitrary accuracy implementation for Gauss-Gegenbauer quadrature based on the iterative method described in [13], which is provided in the software package accompanying this paper.
@@ 2.2 Gauss-Legendre quadrature
With respect to the expansions in terms of elementary functions, the expansion described in [11, sect. 2.2.2], particularizing for $\alpha=\beta=0$, provides an elementary asymptotic expansion for the nodes. An expansion from the weights is obtained by substituting the expansion for the nodes in the expression for the weights in Table 1, considering the expansion for $P_{n}^{\prime}(x)$ for large $n$ and then re-expanding, similarly as was done in [17] for the first few terms. We skip further details of the computation. The results coincide with those obtained in [19] by using the Riemann-Hilbert approach and can be written as follows:
$\begin{array}[]{ll}x_{k}=&c_{k}\left(1-\frac{\displaystyle{1}}{\displaystyle{8}}\kappa^{-2}+\frac{\displaystyle{33+28r_{k}^{2}}}{\displaystyle{384}}\kappa^{-4}-\frac{\displaystyle{865+2060r_{k}^{2}+1208r_{k}^{4}}}{\displaystyle{5120}}\kappa^{-6}+{\cal O}(\kappa^{-8})\right)\\ w_{k}=&\frac{\displaystyle{\pi}}{\displaystyle{\kappa}}s_{k}\left(1-\frac{\displaystyle{1}}{\displaystyle{8}}\kappa^{-2}+\frac{\displaystyle{33+84r_{k}^{2}+56r_{k}^{4}}}{\displaystyle{384}}\kappa^{-4}\right.\\ &\left.-\frac{\displaystyle{865+6180r_{k}^{2}+10160r_{k}^{4}+4832r_{k}^{6}}}{\displaystyle{5120}}\kappa^{-6}+{\cal O}(\kappa^{-8})\right)\end{array}$
@@
Observe that $c_{k}=-c_{n-k+1}$ and $s_{k}=s_{n-k+1}$, as corresponds to the symmetry relations $x_{k}=-x_{n-k+1}$ and $w_{k}=w_{n-k+1}$. This exact symmetry relations are not explicit in the Bogaert formulas, but hold asymptotically.
@@
We have checked that for $n>80$ both expansions (Bessel-type and elementary) coincide within full double precision accuracy for the $36^{th}$ largest node and its corresponding weight, which indicates that the elementary expansions suffice to compute all the nodes and weights except for the largest $35$ positive nodes. This is confirmed by comparing these results with the higher accuracy Maple implementation of Gauss-Gegenbauer quadrature. Additionally, we have checked against the chebfun implementation: our results coincide except that for $n<100$ the accuracy of the chebfun algorithm is slightly worse (when asymptotics is not used). Our algorithm maintains full double precision for any degree.
@@
Our algorithm seems to be faster than those based solely on Bessel-type expansions for large enough $n$, as illustrated in Fig. 1.
Figure 1: CPU times per node and weight spent by the algorithms legpts (chebfun) and legen as a function of the degree $n$ (with $n$ from $10$ to $2000$ with step $10$). The times are computed with Matlab version 2024 running on a laptop with Intel i7-1355U 1.70 GHz processor.

### v2

@@ Abstract
Algorithms for computing the classical Gaussian quadrature rules (Gauss–Jacobi, Gauss–Laguerre, and Gauss–Hermite) are presented, based on globally convergent fourth-order iterative methods combined with asymptotic approximations, which are applied in complementary regions of the parameter space. This approach yields methods that improve upon existing algorithms in speed, accuracy, and computational range. The MATLAB algorithm for Gauss–Jacobi is faster than previous methods and lifts the upper restrictions on the parameters imposed by those methods ($\alpha,\beta\leq 5$); for example, for degrees up to $10^{6}$ all nodes and weights can be computed within the underflow limit for $1<\alpha,\beta\leq 30$, and the computable range of parameters is much larger for smaller degrees, limited only by intrinsic overflow/underflow constraints. For the particular case of Gauss–Legendre quadrature ($\alpha=\beta=0$), a specific asymptotic approach is considered, which yields the most efficient MATLAB implementation available so far. The Gauss–Laguerre and Gauss–Hermite algorithms incorporate subsampling, and scaling is also available in order to extend the computational range. Gauss–Radau and Gauss–Lobatto variants are also considered, along with the computation of the associated barycentric weights. Additionally, arbitrary-precision algorithms (in Maple) are offered for the symmetric cases (Gauss–Gegenbauer and Gauss–Hermite), which can be used to compute thousands of nodes with hundreds of digits in a matter of seconds.
@@ 1 Introduction
where $w(x)$ is a positive weight function defined on the interval $[a,b]$ (with $a$ and/or $b$ possibly infinite), are quadrature rules with maximal polynomial exactness. Specifically, they satisfy $I(x^{k})=Q_{n}(x^{k})$ for $k=0,1,\dots,2n-1$.
As is well known, the nodes $x_{i}$ of an $n$-point Gaussian quadrature rule correspond to the zeros of the orthogonal polynomial $p_{n}(x)$ of degree $n$ (unique up to a normalization factor) that satisfies $I(x^{k}p_{n}(x))=0,\quad k=0,1,\dots,n-1$. The quadrature weights are related to the values of the orthogonal polynomials at these nodes. The nodes and weights can also be expressed in terms of the eigenvalues and eigenvectors of the Jacobi matrix associated with the three-term recurrence relation satisfied by the family of orthogonal polynomials $\{p_{k}(x)\}_{k\in\mathbb{N}}$. See, for instance, [12, sect. 5.3]. Particularly relevant are the so-called classical quadrature rules, for which explicit expressions for the orthogonal polynomials are known. These include Gauss–Jacobi (Gauss–Legendre as a particular case), Gauss–Laguerre and Gauss–Hermite quadratures. In these cases, the weights of the $n$-point rule can be expressed in terms of the derivative of the corresponding orthogonal polynomial evaluated at the nodes. Table 1 summarizes the fundamental relations for these quadratures.
@@
Table 1: Classical Gaussian quadratures: interval of integration $(a,b)$, weight function $w(x)$, orthogonal polynomial $p_{n}(x)$ and its leading order coefficient $k_{n}$ [22, table 18.3.1], and expression for the weights [13]. The constant $M_{n,\alpha,\beta}$ is $M_{n,\alpha,\beta}=2^{\alpha+\beta+1}\Gamma(n+\alpha+1)\Gamma(n+\beta+1)/(n!\Gamma(n+\alpha+\beta+1))$.
Gaussian quadrature rules exhibit fast convergence, particularly for analytic functions [11], making them widely used in numerous applications, including spectral methods. This does not imply, however, the superiority of any Gaussian quadrature rule over other quadrature rules. For instance, it has been observed that the simpler Clenshaw-Curtis quadrature rule is competitive with Gauss–Legendre in terms of convergence rate [31, 32, 34].
On the other hand, the computation of Gauss–Legendre quadratures has dramatically improved, becoming almost elementary in terms of asymptotic approximations [1, 26], as we later explain. Furthermore, for classical Gauss quadratures, several algorithms have been described that can efficiently compute even millions of nodes [1, 2, 13, 14, 15, 17, 20], clearly overruling the well-known Golub-Welsch algorithm [19] (see also [12, sect. 5.3.2]). The algorithms we present in this paper share this capability, while enlarging the range of computation of previous software packages and improving performance. For more general weight functions outside the classical case, these methods are not available, and their computation unavoidably requires slower Golub–Welsch-type approaches (see Gautschi’s monograph [10] and [23] for a more recent example).
Our algorithms for classical Gauss quadratures are based on the asymptotic methods described in [13] (Gauss–Hermite and Gauss–Laguerre) and [15] (Gauss–Jacobi), as well as on the globally convergent iterative methods of [14] (Gauss–Hermite and Gauss–Laguerre) and [17] (Gauss–Jacobi). The MATLAB and Maple programs implementing these algorithms are available in the GitHub repository:
http://github.com/NumericalQuadrature/GaussQuadrature
In comparing our MATLAB algorithms, we consider the most up-to-date MATLAB implementations developed so far, which, in our experience, are the set of algorithms provided with the Chebfun package. The Julia package FastGaussQuadrature.jl [5], largely influenced by Chebfun, is also a worthy contender. Adaptations of our MATLAB algorithms to other programming languages will be considered in the future. Regarding the Maple algorithms, there are no current alternatives, as these are, to our knowledge, the first arbitrary-accuracy algorithms for Gauss–Gegenbauer and Gauss–Hermite quadratures.
The asymptotic and iterative approaches used in our algorithms have distinct advantages and their combination, as we demonstrate in this paper, yields the fastest and most accurate methods available, with the broadest range of validity with respect to the parameters. For example, in the Gauss–Jacobi case our algorithms are not subject to upper restrictions on the parameters $\alpha$ and $\beta$, which both the Chebfun and Julia packages have [^1] The Chebfun algorithm issues a warning when $\max\{\alpha,\beta\}>5$ while the Julia package uses the slower Golub–Welsch algorithm in that case. ($1<\alpha,\beta\leq 5$); the iterative algorithm, being globally convergent and independent of asymptotic approximations, enables the computation for large parameters. The iterative method, however, suffers a mild loss of accuracy as the degree increases, and in that case the purely asymptotic methods are the perfect complement. Similarly, the iterative method allows the computation of Gauss–Laguerre quadratures for almost unbounded values of the parameter $\alpha$, with asymptotic methods as an ideal replacement for large degrees; for Gauss–Laguerre, in addition, scaling can be used to further extend the computational range. Scaling can also be considered in the Gauss–Hermite case. Subsampling is also supported, both for Gauss–Laguerre and Gauss–Hermite quadratures.
@@
The paper is structured as follows. First, for each of the three classical Gaussian quadratures, we describe the different computational methods (iterative and asymptotic), emphasizing aspects specific to each approach that were not fully detailed in previous publications. We begin with Gauss–Jacobi quadrature, with Gauss–Chebyshev and Gauss–Legendre as particularly important cases. For the Legendre case, a highly efficient method based purely on asymptotics and a look-up table for low degrees is presented. After discussing Gauss–Jacobi quadrature, the Gauss–Laguerre and Gauss–Hermite cases are described. Subsequently, we discuss the computation of Gauss–Radau and Gauss–Lobatto quadratures and the evaluation of barycentric weights for interpolation at orthogonal polynomial nodes. Finally, the last section summarizes the set of MATLAB and Maple algorithms accompanying this work.
2 Gauss–Jacobi quadrature
@@
We consider separately the Chebyshev cases $|\alpha|=|\beta|=1/2$, the Legendre case $\alpha=\beta=0$ and the more general Jacobi case (including Gauss–Gegenbauer quadrature $\alpha=\beta$). Before describing these cases in detail, we give a brief summary of some properties of the general Gauss–Jacobi case that will be useful in the subsequent discussion.
@@
The values $|\dot{Y}(\theta_{i})|$ have a small variation as a function of $i$. This is a consequence of the fact that the coefficient of the ODE (2) is nearly constant as $n\rightarrow+\infty$ in any compact subinterval of (1,1)(-1,1) (and therefore the solutions are close to trigonometric functions in a first approximation). This is consistent with the known asymptotic approximations as $n\rightarrow+\infty$ (see for instance [26], Eq. (A.8)):
@@
2.1 Gauss–Chebyshev quadratures
For the Chebyshev cases, $|\alpha|=|\beta|=1/2$, simple explicit formulas exist for the nodes and weights [25, Thm. 8.4]. The four cases can be summarized in a single formula as follows:
@@ 2.1 Gauss–Chebyshev quadratures
Care must be taken in order to avoid loss of accuracy in the evaluation of the trigonometric functions involved in (6). For instance, the previous expression for $x_{k}$ is not convenient for computing the nodes close to zero which, if $\alpha<<n$, correspond to values $k\simeq n/2$. Indeed, we have $x_{k}=\cos\theta_{k}$ with $\theta_{k}=\pi/2+\delta_{k}$, with $|\delta_{k}|$ small; then $x_{k}=-\sin(\delta_{k})\simeq-\delta_{k}$ and $\delta_{k}=\theta_{k}-\pi/2$ suffers loss of accuracy by cancellation. In this case it is preferable to use the expression $x_{k}=-\sin((\frac{n}{2}-k+\frac{\alpha-\beta+2}{4})\frac{\displaystyle{\pi}}{\displaystyle{\kappa}})$ where the small number n/2kn/2-k (compared to $\kappa$) is computed exactly. Similarly precautions must be taken for the factors $1\pm x_{k}$ appearing in the weights.
@@
2.2 Gauss–Legendre quadrature
The Gauss–Legendre case is a particularly important one and, in fact, it is customary to use the expression “Gauss quadrature” to refer to this method alone. Because $x_{k}=-x_{n-k+1}$ ($1<x_{1}<...<x_{n}<1$) and $w_{k}=w_{n-k+1}$ we only need to compute the non-negative nodes and its corresponding weights.
For Gauss–Legendre quadrature, the Bessel-type expansions of Legendre polynomials [28] enable explicit asymptotic expressions for nodes and weights, as in [1]. The Chebfun package legpts [4] follows this approach for $n>100$. We also use Bogaert’s expansions, but only for a fixed number of nodes and weights: for $n>80$, the 35 largest positive nodes and corresponding weights $w_{i}$ are computed using Bogaert’s formulas, while expansions in elementary functions are considered for the rest. For $n\leq 80$, we use a look-up table storing precomputed nodes and weights with full double precision ($16$ significant digits), generated using the arbitrary accuracy implementation for Gauss–Gegenbauer quadrature based on the iterative method described in [17], which is provided in the software package accompanying this paper.
@@ 2.2 Gauss–Legendre quadrature
With respect to the expansions in terms of elementary functions, the expansion described in [15, sect. 2.2.2], particularizing for $\alpha=\beta=0$, provides an elementary asymptotic approximation for the nodes. An expansion for the weights is obtained by substituting the expansion for the nodes in the expression for the weights in Table 1, considering the expansion for $P_{n}^{\prime}(x)$ for large $n$ and then re-expanding, similarly as was done in [24] for the first few terms. We skip further details of the computation. The results coincide with those obtained in [26] by using the Riemann-Hilbert approach and can be written as follows:
(9) $\begin{array}[]{ll}x_{k}=&c_{k}\left(1-\frac{\displaystyle{1}}{\displaystyle{8\kappa^{2}}}+\frac{\displaystyle{33+28r_{k}^{2}}}{\displaystyle{384\kappa^{4}}}-\frac{\displaystyle{865+2060r_{k}^{2}+1208r_{k}^{4}}}{\displaystyle{5120\kappa^{6}}}+{\cal O}(\kappa^{-8})\right)\\ &\\ w_{k}=&\frac{\displaystyle{\pi}}{\displaystyle{\kappa}}s_{k}\left(1-\frac{\displaystyle{1}}{\displaystyle{8\kappa^{2}}}+\frac{\displaystyle{33+84r_{k}^{2}+56r_{k}^{4}}}{\displaystyle{384\kappa^{4}}}\right.\\ &\\ &\left.-\frac{\displaystyle{865+6180r_{k}^{2}+10160r_{k}^{4}+4832r_{k}^{6}}}{\displaystyle{5120\kappa^{6}}}+{\cal O}(\kappa^{-8})\right)\end{array}$
@@
Observe that $c_{k}=-c_{n-k+1}$ and $s_{k}=s_{n-k+1}$, as corresponds to the symmetry relations $x_{k}=-x_{n-k+1}$ and $w_{k}=w_{n-k+1}$. The exact symmetry relations are not explicit in the Bogaert formulas, but hold asymptotically.
@@
For $n>80$, both expansions (Bessel-type and elementary) coincide within full double precision accuracy for the $36^{th}$ largest node and its corresponding weight; this indicates that the elementary expansions suffice to compute all the nodes and weights except for the largest $35$ positive nodes. This is confirmed by comparing these results with the higher accuracy Maple implementation of Gauss–Gegenbauer quadrature. Additionally, we have compared against the Chebfun implementation: our results coincide except that for $n<100$ the accuracy of the Chebfun algorithm is slightly worse (when asymptotics is not used). Our MATLAB algorithm maintains full double precision for any degree, as the comparison against the Maple implementation shows.
Summarizing, we consider the following scheme for computing Gauss–Legendre quadratures:
  1. 1.
    Compute the non-negative nodes $x_{k}$, $k=\lfloor n/2\rfloor+1\ldots n$ and the corresponding weights $w_{k}$:
    1. (a)
      If $n\leq 80$ use the look-up table.
    2. (b)
      If $n>80$ use asymptotic approximations as follows:
      1. i.
        For the largest $35$ positive nodes and its corresponding weights ($x_{k}$, $w_{k}$, $k=n-34,\ldots n$) use Bogaert’s formulas (7) and (8).
      2. ii.
        Compute the rest of non-negative nodes and corresponding weights ($x_{k}$, $w_{k}$, $k=\lfloor\frac{n}{2}\rfloor+1\ldots n-35$) with the elementary expansions (9).
  2. 2.
    Use symmetry for the negative nodes and the corresponding weights: $x_{k}=-x_{n-k+1}$ $w_{k}=w_{n-k+1}$, $k=1,\ldots\lfloor n/2\rfloor$
@@
Our algorithm seems to be faster than those based solely on Bessel-type expansions for large enough $n$, as illustrated in Fig. 1. In this figure, the times are measured with the MATLAB instruction timeit. Both codes are so simple and efficient that the time spent is too small for a reliable measurement, and we have averaged the time over $100$ cycles of computation for each node. Needless to say, the detail of the comparative result will depend on the performance of the computer; the results in Fig. 1 correspond to execution times on a typical laptop, specifically a laptop equipped with and Intel(R) Core(TM) i7-1355U processor with 10 cores (12 threads). The difference can be smaller in higher performance computers. For instance, in an Intel(R) Core(TM) i7-14700K with 20 cores (28 threads), we observe roughly a reduction by a factor $2$ in the relative difference (not shown), but legen is still more that $5$ times faster than legpts for $n\leq 2000$.
Figure 1: CPU times in seconds per node and weight spent by the algorithms legpts (Chebfun) and legen as a function of the degree $n$ (with $n$ from $10$ to $2000$ with step $10$). The times are computed with MATLAB version 2024b running on a laptop with Intel(R) Core(TM) i7-1355U processor. The times are averaged over $100$ evaluations for each value of $n$.